OpenStudy (anonymous):

Solve the proportion. 16/50=x/156.25 @Destinymasha

OpenStudy (texaschic101):

16/50 = x/156.25 cross multiply (50)(x) = (16)(156.25) 50x = 2500 x = 2500/50 x = 50 check... 16/50 = 50/156.25 cross multiply 50(50) = (16)(156.25) 2500 = 2500 (correct) x = 50 any questions ?

OpenStudy (anonymous):

waitu worked it out with the answer?

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

the answer is : x = 50 I found the answer then checked it by subbing it back into the proportion